OpenStudy (anonymous):

Solve : 2/5 (x − 4) = 2x for x.

OpenStudy (anonymous):

2/5(x-4)=2x (5/2)(2/5)(x-4)=(2x)(5/2) (x-4)=5x -4=5x-x -4=4x -1=x

OpenStudy (anonymous):

Do I divide 2 by 5 before doing distributive?

OpenStudy (strawberry17):

No, you can just multiply 2/5 by x and -4. I think it's easier..

OpenStudy (anonymous):

i don't like fraction so for me it's easier changing it too decimal that it would be: 0.4x-1.6=2x -1.6=2x-0.4x -1.6=1.6x -1.6/1.6=1.6x/1.6 -1=x

OpenStudy (strawberry17):

ya, that would work too..

OpenStudy (anonymous):

what i did was got rid of the fraction for you, so you dont have to deal with fractions or decimals, but its up to you :)

OpenStudy (strawberry17):

okay

OpenStudy (anonymous):

either one you choose it will give you the same answer just different ways just choose the one that's easier for you to use and remember it's just about letting the variable alone